OpenStudy (anonymous):

While solving (3-8i)^2-(3+8i)^2 should i solve for the exponents first? or FOIL?

OpenStudy (amistre64):

the only thing that will get you further is to ... foil... as you call it

OpenStudy (amistre64):

of course now that i read it again, it comes across a bit vague; wht do yo mean by takeing care of exponents?

OpenStudy (amistre64):

(3-8i)^2 , expand this out -(3+8i)^2 ; expand this out and dont forget to apply the negative

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

im still unsure ive expanded the prblem and solved for it but didnt get the correct answer

OpenStudy (anonymous):

Do: (3-8i)(3-8i) first. Then do (3+8i)(3+8i) So, you should have [9-24i-24i+64i^2] minus [9+24i+24i+64i^2]

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

You're welcome. hope it helped

OpenStudy (anonymous):

indeed

OpenStudy (anonymous):

final answer 96i

OpenStudy (anonymous):

-96i

OpenStudy (anonymous):

Yes! correct. This site will be your best friend. http://www.wolframalpha.com/input/?i=%283-8i%29^2-%283%2B8i%29^2+ Sometimes it even shows you the steps.

OpenStudy (anonymous):

I think the most direct method is to factor by difference of squares and simplify.\[(3-8i)^2-(3+8i)^2=(3-8i+3+8i)(3-8i-3-8i)=6(-16i)=-96i\]

OpenStudy (amistre64):

good eye ;)

OpenStudy (anonymous):

Nice to get one correct occasionally.